### v1

# Augmentation-based Unsupervised Cross-Domain Functional MRI Adaptation for Major Depressive Disorder Identification

Abstract
Major depressive disorder (MDD) is a common mental disorder that typically affects a person’s mood, cognition, behavior, and physical health. Resting-state functional magnetic resonance imaging (rs-fMRI) data are widely used for computer-aided diagnosis of MDD. While multi-site fMRI data can provide more data for training reliable diagnostic models, significant cross-site data heterogeneity would result in poor model generalizability. Many domain adaptation methods are designed to reduce the distributional differences between sites to some extent, but usually ignore overfitting problem of the model on the source domain. Intuitively, target data augmentation can alleviate the overfitting problem by forcing the model to learn more generalized features and reduce the dependence on source domain data. In this work, we propose a new augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework for automatic diagnosis of MDD. The AUFA consists of 1) a graph representation learning module for extracting rs-fMRI features with spatial attention, 2) a domain adaptation module for feature alignment between source and target data, 3) an augmentation-based self-optimization module for alleviating model overfitting on the source domain, and 4) a classification module. Experimental results on 1,089 subjects suggest that AUFA outperforms several state-of-the-art methods in MDD identification. Our approach not only reduces data heterogeneity between different sites, but also localizes disease-related functional connectivity abnormalities and provides interpretability for the model.
Index Terms:
Resting-state functional MRI; Domain adaptation; Multi-site data; Major depressive disorder
I Introduction
Figure 1: Overview of our augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework, including (1) fMRI feature extraction for brain functional connectivity networks using a Transformer module, (2) domain adaptation module for feature alignment between source and target domains, (3) augmentation-based self-optimization module, and (4) a classification module. The target data are fed into the model that has been jointly optimized by $L_{C}$, $L_{M}$ and $L_{A}$ for direct prediction.
Major depressive disorder (MDD) is one of the most common mental health disorders, leading to symptoms such as extremely low mood, self-denial, feelings of helplessness, and impaired cognitive function, thus significantly impacting the patient’s daily life and work [1]. Currently, the diagnosis of MDD mainly relies on the patient’s symptom assessment, physical examination, and psychological scales, which still involves subjectivity and risk of misdiagnosis [2, 3]. Therefore, accurate detection of MDD is crucial in helping clinicians to formulate timely intervention plans.
Resting-state functional magnetic resonance imaging (rs-fMRI), as a non-invasive and non-radioactive technique, can assess changes in neural activity across the brain over time. It has been widely used in computer-aided analysis of brain disorders using learning-based methods [4, 5]. Currently, machine learning/deep learning techniques generally focus on using functional connectivity networks (FCN) derived from blood-oxygen-level dependent (BOLD) time series of rs-fMRI data to capture interactions between brain regions-of-interest (ROIs) [6]. Traditional machine learning methods often extract handcrafted features based on FCN for brain disease detection. However, these methods require prior knowledge of FCN [7], and shallow features usually result in poor performance in classification tasks.
With the development of deep learning, numerous methods employ graph neural networks to automatically learn high-level feature representations of fMRI data for automatic diagnosis of MDD [8, 9]. For example, Li et al[10] extracted topological features from brain functional connectivity constructed based on fMRI data, and utilized a graph neural network (GNN) to identify MDD. Jun et al[11] combined valid connectivity features from rs-fMRI with non-imaging phenotypic features, and the combined features were further learned by a graph convolutional network (GCN). However, these methods only use a single site from the MDD dataset for experiments, which does not guarantee the generalization ability of the model. Multi-site rs-fMRI data provide more data for learning-based methods to train reliable models [12, 13]. However, fMRI acquired from different sites often exhibit significant heterogeneity due to equipment differences and variations in data acquisition conditions [14], which presents a challenge for diagnosing brain diseases.
Domain adaptation helps reduce the data distribution difference between the source and target domains [15]. However, when the gap between source and target domains is large, coupled with the size limitation of fMRI datasets, it becomes easier for the model to overfit on the source domain  [16]. That is, the model is too fine-grained in capturing the source domain features and ignores the generalization of the universal laws. Intuitively, augmenting the target domain data can enhance the model’s generalization ability to the target domain and avoid over-reliance on specific source domain data [17, 18].
To this end, we propose a new augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework for automated diagnosis of MDD. As shown in Fig. 1, we first construct FCNs for source/target domain subjects based on fMRI time series. Then, the Transformer encoder is utilized to learn fMRI features with spatial attention. At the same time, we introduce the maximum mean difference constraint to align the fMRI features learned from the source and target domains, aiming for the Transformer encoder to learn the shared features between domains. In addition, to further improve the model’s generalization ability to the target domain, we optimize the model by augmenting the data in the target domain to prevent the model from overfitting on the source domain. AUFA can significantly reduce the impact of data distribution differences across different sites through the supervision of source domain labels and adaptation of data between domains. Experiments on rs-fMRI data from 1,089 subjects demonstrate the excellent performance of AUFA in MDD identification compared with competing methods.
The major contributions of this work are summarized below.
  • A novel augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework is proposed for automated diagnosis with resting-state fMRI data.
  • The proposed AUFA introduces the maximum mean difference constraint to align fMRI features from different sites, which can effectively reduce the data distribution differences between source and target domains.
  • The self-optimization module based on augmentation in AUFA helps prevent the overfitting of the model in the source domain and improves the generalization ability in the target domain.
  • AUFA can effectively capture the attention relationship between different brain regions to automatically learn fMRI features and obtain discriminative functional connectivity.
  • Cross-domain disease detection experiments at 5 fMRI imaging sites show that our method is superior to several state-of-the-art methods.
The remainder of this paper is organized as follows. Section II provides a brief review of related studies. Section III introduces the data acquisition and pre-processing. Section IV introduces the details of the proposed method. In Section V, we present the experimental setup, competing methods, and results. Section VI discusses key components of the proposed method and list several limitations of the current work. Finally, we summarize the paper in Section VII.
II Related Work
II-A Learning-based Methods for Functional MRI Analysis
Deep learning, as a powerful machine learning method, has been widely used in fMRI analysis [19, 20]. Compared with traditional machine learning methods, deep learning models can better handle high-dimensional, nonlinear and complex fMRI data with better modeling capabilities and generalization performance [21]. Numerous studies have used multilayer perceptron or convolutional neural network models to extract features of FCNs constructed based on fMRI data [22]. Graph neural networks, as a class of deep learning algorithms dedicated to processing graph data, have also been widely used for the analysis of FCN based on fMRI, such as graph convolutional networks (GCNs) [23]. GCNs learn node representations by aggregating features of nodes and their neighbors through convolutional operations. Yao et al.et al[24] proposed a mutual multi-scale triplet GCN to analyze brain functional and structural connectivity for brain disease diagnosis. Noman et al.et al[25] designed a novel graph autoencoder (GAE) architecture based on GCN for MDD detection. However, these methods usually overlook the different contributions of different brain regions to the classification task. Transformer is an attention-based deep learning model originally proposed for natural language processing tasks such as machine translation [26]. It introduces a self-attention mechanism that enables modeling the dependencies between different positions in an input sequence. In recent years, the Transformer model has also been applied as a powerful sequence modeling tool for capturing attention relationships between brain regions in fMRI data analysis [27, 28]. In this work, we attempt to leverage the Transformer in unsupervised domain adaptation for multi-site fMRI feature extraction. Considering that the strong backbone of Transformer increases the risk of overfitting, we augment the target domain data and improve the model generalization over the target domain through a self-optimization module.
II-B Domain Adaptation of Functional MRI Data
The multi-site fMRI data greatly increased the sample size for the model training. However, due to distributional differences among data from different sites, the training data may not accurately reflect the distribution of the test data, leading to a degradation in model performance [29, 14]. The domain adaptation approach is an important strategy for solving this problem. Currently, domain adaptation methods can be classified into three categories based on the availability of target domain labels [30]: (1) supervised domain adaptation, where a small amount of labeled data in the target domain can be used for model training [31]; (2) semi-supervised domain adaptation, where a small amount of labeled data and redundant unlabeled data in the target domain are available during the training process [32]; and (3) unsupervised domain adaptation, where only unlabeled target data can be used to train the adaptive model [33]. In this paper, we focus on unsupervised domain adaptation.
In recent years, various domain adaptation methods have been widely used in the field of computer vision [34]. For example, Zhou et al[35] proposed a new Adversarial Distribution Adaptation Network (ADAN) to jointly reduce the global and local distributional differences between different domains to learn domain invariant representations. Li et al[36] proposed a Progressive Adaptation of Subspace (PAS) method to gradually obtain reliable pseudo-labels for model training. Liu et al[37] developed a novel unsupervised domain adaptation framework that ensures the preservation of the source data distribution to guide target structure learning in a semi-supervised clustering manner. However, multi-site adaptation studies for fMRI time series are still limited. Some studies based on multi-site fMRI data have simply combined fMRI data from different sites for training [38], i.e., assuming that they have the same data distribution, ignoring data heterogeneity between sites. Even though some studies attempt to reduce the distributional differences of fMRI data through domain adaptation methods, they usually ignore the problem that the model is likely to be overfitted on the source domain. For example, Wang et al[39] proposed an unsupervised graph domain adaptation network for the diagnosis of autism spectrum disorders, which consists of three main components: graph isomorphism encoder, progressive feature alignment, and unsupervised information matrix regularization. Lee et al[29] designed a new domain generalization meta-learning model for the diagnosis of brain diseases based on rs-fMRI. The aim of this paper is to study unsupervised domain adaptation for multi-site fMRI data from two perspectives, i.e., reducing inter-domain distributional differences and mitigating the risk of model overfitting.
TABLE I: Phenotypic information of 1089 subjects from REST-meta-MDD consortium [40]. MDD: Major depressive disorder; NC: Normal control; M/F: Male/Female. The age information is reported as mean±$\pm$±standard deviation.
Site ID Category Subject Gender (M/F) Age (Years)
      20 MDD 282 $99/183$ $38.74\pm 13.63$
NC 251 $87/164$ $39.64\pm 15.84$
      21 MDD 86 $38/48$ $34.71\pm 12.56$
NC 70 $31/39$ $36.13\pm 12.55$
      25 MDD 89 $21/68$ $65.60\pm 6.72$
NC 63 $29/34$ $69.63\pm 5.82$
       1 MDD 74 $31/43$ $31.72\pm 8.14$
NC 74 $32/42$ $31.80\pm 8.93$
       9 MDD 50 $23/27$ $28.10\pm 8.74$
NC 50 $31/19$ $28.92\pm 8.50$
III Materials and Data Preprocessing
III-A Data Acquisition
This work involved rs-fMRI data from the top five largest sites in the imaging center of the REST-meta-MDD consortium [40], including SITE 20, SITE 21, SITE 25, SITE 1, and SITE 9. There were a total of 1,089 participants across the five sites, including 581 MDD patients and 508 normal controls (NCs). The rs-fMRI data from SITE 20, SITE 21 and SITE 1 were acquired by Siemens Verio 3T with the following parameters: echo time (TE) $=30\,ms$, repetition time (TR) $=2,000\,ms$, flip angle $=90$. Particularly, at SITE 20, the thickness/gap was $3.0\,mm/1.0\,mm$ with $32$ slice; at SITE 21, the thickness/gap was $3.5\,mm/0.7\,mm$ with $32$ slice, and at SITE 1, the thickness/gap was $4.0\,mm/0.8\,mm$ with $32$ slice. The rs-fMRI data from SITE 25 were obtained by Siemens Verio 3T with the following parameters: TE $=25\,ms$, TR $=2,000\,ms$, flip angle$=90$, thickness/gap $=4.0\,mm/0\,mm$, slice number $=36$. The rs-fMRI data for SITE 9 were obtained by GE Discovery MR750 3.0T with the following parameters: TE $=25\,ms$, TR $=2,000\,ms$, flip angle$=90$, thickness/gap = $3.0\,mm/1.0\,mm$, slice number $=35$. More detailed demographic information is reported in Table I.
Figure 2: Graph representation learning module in AUFA that learns new features with spatial attention through multiple Transformer encoder layers.
III-B Image Pre-Processing
We pre-processed the initial rs-fMRI data through the Data Processing Assistant for Resting-State fMRI (DPARSF) [41] pipeline, including the following steps: (1) discard the first ten volumes, (2) head motion correction, (3) space normalization and smoothing, (4) interference signal removal, (5) bandpass filtering. Finally, we extracted an average rs-fMRI time series for 116 predefined regions of interest (ROI) based on Automated Anatomical Labeling (AAL) [42] atlas.
IV Proposed Method
IV-A Problem Formulation
In this work, we focus on using cross-domain rs-fMRI data for the MDD identification task, aiming to solve an interesting, realistic, but challenging unsupervised domain adaptation problem, i.e., utilizing information learned from the source domain to assist the classification of a completely unlabeled target domain. We first introduce some notation that will be used in this work. Let $X^{d}$ be the input (graphs constructed based on fMRI data) space and $Y^{d}$ be the target (label) space, where $d$ identifies either the source domain $s$ or the target domain $t$. The source domain is defined as $S$ = $\{X^{s},Y^{s}\}$, containing the feature set $X^{s}$ and the corresponding label set $Y^{s}$. The target domain is defined accordingly as $T$ = $\{X^{t},Y^{t}\}$ with feature set $X^{t}$ and label set $Y^{t}$. Even though the two domains have the same task (they share the same label space $Y^{s}$ = $Y^{t}$), their data distribution is somewhat different, i.e., $P^{s}(X)\neq P^{t}(X)$. In the framework of unsupervised domain adaptation, the challenge is how to learn a feature extractor $\phi:X\rightarrow Z$ that leverages the abundant labeling information available in the source domain. This extractor maps the initial input to a feature space $Z$, aiming to minimize the data distribution between the source and target domains, i.e., $P^{s}(\phi(X))\approx P^{t}(\phi(X))$.
IV-B Method
We propose an augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework, which is trained using labeled source domain data and unlabeled target domain data for prediction in the target domain. As shown in Fig. 1, the AUFA comprises four main components: (1) graph representation learning for extracting rs-fMRI features based on Transformer from the constructed brain FCNs, (2) domain adaptation for feature alignment, (3) augmentation-based self-optimization module, and (4) classification.
IV-B1 Graph Representation Learning Module
This module includes two parts: graph construction and Transformer-based graph representation learning, with details introduced below.
Graph Construction. The brain is a complex neural network system, wherein different ROIs are responsible for respective functions and interact with each other to coordinate cognitive processes and behaviors. The FCN constructed by analyzing the fMRI time series relationship between different ROIs can effectively capture the complex information of the brain. Each subject’s FCN can be modeled as an undirected graph $G$ = $(V,E)$, where $V$ is the node set, representing different ROIs, and $E$ is the edge set, representing the strength of functional connections between ROI pairs. Since FCN can capture the topology of the brain, it is widely used for computer-aided diagnosis of brain diseases.
In this work, we first construct FCNs for subjects from the source domain and the target domain, respectively. Specifically, we use Pearson correlation (PC) to calculate correlations of rs-fMRI time series, establishing functional connections between ROI pairs. The functional connection $w_{ij}\in\ [-1,1]$ between the $i$-th ROI and the $j$-th ROI is defined as:
$w^{d}_{ij}=\frac{(x^{d}_{i}-\bar{x}^{d}_{i})^{T}(x^{d}_{j}-\bar{x}^{d}_{j})}{\sqrt{(x^{d}_{i}-\bar{x}^{d}_{i})^{T}(x^{d}_{i}-\bar{x}^{d}_{i})}\sqrt{(x^{d}_{j}-\bar{x}^{d}_{j})^{T}(x^{d}_{j}-\bar{x}^{d}_{j})}},$ (1)
where $\bar{x}^{d}_{i}$ denotes the mean vector of the BOLD signal ${x}^{d}_{i}$ corresponding to the $i$-th ROI, with $d$ representing the source or target domain. Thus, for each subject, we obtain the functional connectivity matrix $X^{d}$ , where $X^{d}_{ij}$=$w^{d}_{ij}$, and utilize the constructed $X^{d}\in R^{N\times N}$ as the initial graph feature for the source and target domains.
Transformer-based Graph Representation Learning. To accurately capture the specific contributions of individual brain regions and learn more discriminative features, we adopted the Transformer encoder for graph representation learning. This strategy utilizes the Transformer’s self-attention mechanism to weigh the importance of different brain regions and obtain the corresponding attention coefficients. Subsequently, based on the attention weights, the feature information of global nodes is aggregated and updated. Thus, it facilitates a deeper understanding of the structure of complex brain networks. More details are introduced in the following.
We take the node features $X^{d}$ as input, and model the relationship between each node and all other nodes through the self-attention mechanism. The attention scores between nodes are calculated and the corresponding node representations are weighted and aggregated. As shown in Fig. 2, define $Z^{l}(Z^{1}=X^{d})$ to be the graph representation of some subject in layer $l$. Taking the self-attention module in the $h$-th header as an example, this component is computed as follows:
$Z^{h,l}$ $=Attention(Z^{l}W^{h,l}_{Q},Z^{l}W^{h,l}_{K},Z^{l}W^{h,l}_{V})=A^{h,l}Z^{l}W^{h,l}_{V},$ (2)
where $A^{h,l}=softmax(\frac{Z^{l}W^{h,l}_{Q}(Z^{l}W^{h,l}_{K})^{T}}{\sqrt{D_{k}}})\in R^{N\times N}$ is the matrix of attention scores among all nodes in the $l$-th layer of the $h$-th header, $W^{h,l}_{Q}$, $W^{h,l}_{K}$, $W^{h,l}_{V}$ is the matrix of learnable linear transformations exerted on the node features, and $D_{k}$ is the dimension of $Z^{l}W^{h,l}_{K}$.
To capture more comprehensive information about the graph data, we use a multi-head self-attention mechanism for parallel computation. The input $X^{d}$ is passed to $H$ different self-attention modules and $H$ output matrices $Z$ are computed and aggregated. The aggregation operation can be expressed as:
$\bar{Z^{l}}=N(W(C(Z^{1,l},Z^{2,l},...,Z^{H,l}))),$ (3)
where $C$ is the concatenation operation, $W$ denotes linear transform, and $N$ denotes layer normalization.
Based on the output of the self-attention layer, a further nonlinear transformation is performed by a feed-forward neural network. The feed-forward neural network is a two-layer fully connected neural network with the following formulation:
${Z^{l+1}}=N(\Psi(\bar{Z}^{l}W_{1}+b_{1})W_{2}+b_{2}),$ (4)
where $\Psi$ denotes the $RELU$ activation function and $W_{1}$, $W_{2}$, $b_{1}$, $b_{2}$ are the learnable weights and biases. Thus, a new feature representation $Z_{X^{d}}$ of $X^{d}$ is obtained through the learning of multiple Transformer layers. To conduct the subsequent graph classification task, the $Z_{X^{d}}$ are flattened to obtain the final graph-level representation $f_{X^{d}}$.
IV-B2 Domain Adaptation Module
To reduce domain offset, we employ a maximum mean difference constraint to align the learned feature on different domains. As shown in Fig. 1, we obtain the graph-level representation $f_{X^{d}}$ of the source and target domains through the Transformer module learning. The $f_{X^{d}}$ are then fed into two fully connected layers, which are finally predicted by a softmax layer.
Here, considering that the low-level features contain limited semantic information, we only align the graph-level feature mappings on the source and target domains, represented as follows:
$L_{M}=\frac{1}{B^{2}}\sum_{l=1}^{L}\sum_{i=1}^{n}\sum_{j=1}^{n}(f_{s}^{l(i)}-f_{t}^{l(j)})^{T}(f_{s}^{l(i)}-f_{t}^{l(j)}),$ (5)
where $f_{s}^{l(i)}$ is the features of the $i$-th sample in the source domains, $f_{t}^{l(j)}$ are the features of the $j$-th sample in the target domains, $B$ is the batch size, $n$ is the number of samples, and $L$ denotes the number of layers in the fully connected layer. Eq. (5) measures the discrepancy in feature distribution between the two domains. By minimizing this equation, we penalize the features with different distributions across domains, compelling the Transformer feature extractor to learn shareable feature representations across the two domains.
IV-B3 Augmentation-based Self-optimization Module
In Section IV-B1, we utilize a Transformer encoder to learn the graph representation for each subject. However, due to the limited training data and the large number of parameters in the multilayered self-attention mechanism and the feed-forward neural network, the model is prone to overfitting. In the unsupervised domain adaptation setting, the risk of the model overfitting the source domain data is significantly heightened as a result of the model’s reliance solely on supervision from the source domain labels. Therefore, we reduce the risk of model overfitting on the source domain data by augmenting the target domain data to improve model generalizability.
To enhance model robustness, we randomly select the input features of the $l$-th layer Transformer encoder for augmentation in the target domain. Given the initial graph feature $X^{t}$ to be augmented, we utilize another randomly selected subject $X^{t^{\prime}}$ from the target domain to add an offset to $X^{t}$ in order to achieve augmentation. Let $Z^{l}_{X^{t}}$, $Z^{l}_{X^{t^{\prime}}}$ denote the input features of $X^{t}$ and $X^{t^{\prime}}$ in layer $l$, respectively. The features of subject $X^{a}$ in the $l$-th layer after augmentation can be expressed as:
$Z^{l}_{X^{a}}=Z^{l}_{X^{t}}+\alpha(Z^{l}_{X^{t^{\prime}}}-{Z}^{l}_{X^{t}}),$ (6)
where $\alpha$ is a scalar that is used to adjust the degree of offset to the target domain features $X^{t}$. In fact, such an augmentation operation is equivalent to imposing regularization constraints on the model. It is worth noting that by randomly selecting the $l$-th layer for augmentation, the learning of subsequent parameters from the $(l+1)$-th layer to the last layer will be affected, thus achieving a multilayer regularization constraint on the model. As shown in Fig. 1, $Z^{l}_{X^{t}}$ is augmented to the new target domain data $Z^{l}_{X^{a}}$ after the perturbation of $Z^{l}_{X^{t^{\prime}}}$ in the $l$-th layer. Then, the new feature representations $Z_{X^{t}}$ and $Z_{X^{a}}$ are obtained through the subsequent learning of the Transformer layer. We calculate the prediction probabilities of $Z_{X^{t}}$ and $Z_{X^{a}}$ by performing the same procedure as mentioned (i.e., flatten, fully connected layers and softmax), and denoted as $P_{X^{t}}$, $P_{X^{a}}$. Then, we compute the Kullback-Leibler ($KL$) divergence between $P_{X^{t}}$ and $P_{X^{a}}$ to minimize it:
$KL(P_{i}||P_{j})=\sum P_{i}~{}log\frac{P_{i}}{P_{j}},$ (7)
where $i,j\in\{X^{t},X^{a}\},i\neq j$. We update the model parameters according to the joint effect of $P_{X^{t}}$ and $P_{X^{a}}$ to avoid excessive gradients generated by any single probability.
It is worth noting that we do not have label information for the target domain to determine whether predictions about the target domain are reliable. Therefore, a confidence filter is added to round off samples corresponding to prediction probabilities with low confidence. The set of remaining target domain data can be represented as:
$F=\{x\in D|max(P_{i})\textgreater\epsilon\},$ (8)
where $D=\{X^{t}\cup X^{a}\}$ and $\epsilon$ is a threshold for filtering the prediction probability. Thus, the final augmentation-based self-optimization loss is defined as:
$L_{A}=KL(F(P_{i})||P_{j}),$ (9)
where $i,j\in\{X^{t},X^{a}\}$, $i\neq j$.
IV-B4 Classification
The resulting graph-level representation of each subject is fed into two fully connected layers for classification, and the prediction probability vector is obtained through the softmax layer. The number of neurons in the first and second layers was $4096$ and $2$, respectively. The whole model is trained using a joint loss function $L$ as the optimization objective, which comprises three components: (1) cross-entropy loss $L_{C}$ measures the difference between the model’s prediction for source domain data and the real label to ensure the model’s classification performance in source domain, (2) maximum average difference loss $L_{M}$ is used to align the learned inter-domain features, (3) augmentation-based self-optimization loss $L_{A}$ is employed to further optimize the model and prevent overfitting of the model on the source domain data. Therefore, the joint loss function can be expressed as:
$L=L_{C}+\lambda_{1}L_{M}+\lambda_{2}L_{A},$ (10)
where $\lambda_{1}$, $\lambda_{2}$ are hyperparameters.
IV-C Implementation Details
The proposed AUFA model is implemented on Pytorch using NVIDIA Quadro RTX 6000. The model is trained in two stages. In the first stage, the AUFA is trained in a supervised manner based on the source domain to narrow down the parameter search space. In the second stage, the pre-trained AUFA is further optimized by minimizing the joint loss function Eq. (10) based on all labeled source domain data and unlabeled target domain data. We set the number of layers in the Transformer encoder to $2$ and the number of attention heads to $4$. Hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are both $1$. The AUFA model is optimized by Adam algorithm with a learning rate of $0.00001$, training epoch of $45$ and batch size of $32$.
TABLE II: Results (mean±$\pm$±standard deviation) of eight methods in MDD vs. NC classification, and the best results are shown in bold.
Method Accuracy Precision Recall AUC F1-score Accuracy Precision Recall AUC F1-score
SITE 20 $\rightarrow$ SITE 21 SITE 20 $\rightarrow$ SITE 25
BC-SVM 0.51 0.55 0.63 0.49 0.59 0.51 0.57 0.66 0.49 0.61
LE-SVM 0.52 0.56 0.64 0.52 0.59 0.53 0.61 0.57 0.50 0.59
MNSFS 0.48 0.56 0.29 0.49 0.38 0.49 0.58 0.45 0.52 0.51
GIN 0.52$\pm$0.04 0.57$\pm$0.03 0.55$\pm$0.13 0.51$\pm$0.03 0.55$\pm$0.08 0.50$\pm$0.04 0.60$\pm$0.05 0.50$\pm$0.12 0.48$\pm$0.06 0.53$\pm$0.07
ST-GCN 0.52$\pm$0.05 0.56$\pm$0.04 0.60$\pm$0.18 0.49$\pm$0.04 0.57$\pm$0.08 0.57$\pm$0.01 0.62$\pm$0.02 0.71$\pm$0.09 0.55$\pm$0.03 0.65$\pm$0.03
CORAL 0.54$\pm$0.02 0.59$\pm$0.01 0.58$\pm$0.02 0.56$\pm$0.03 0.59$\pm$0.01 0.57$\pm$0.02 0.67$\pm$0.04 0.56$\pm$0.01 0.60$\pm$0.03 0.61$\pm$0.02
DANN 0.56$\pm$0.03 0.62$\pm$0.06 0.56$\pm$0.13 0.58$\pm$0.06 0.57$\pm$0.04 0.54$\pm$0.09 0.60$\pm$0.07 0.58$\pm$0.28 0.55$\pm$0.08 0.57$\pm$0.15
AUFA (Ours) 0.63$\pm$0.03 0.66$\pm$0.03 0.73$\pm$0.06 0.62$\pm$0.04 0.69$\pm$0.03 0.60$\pm$0.04 0.68$\pm$0.04 0.64$\pm$0.03 0.59$\pm$0.04 0.66$\pm$0.03
SITE 20 $\rightarrow$ SITE 1 SITE 20 $\rightarrow$ SITE 9
BC-SVM 0.49 0.49 0.58 0.49 0.53 0.51 0.51 0.74 0.46 0.60
LE-SVM 0.51 0.52 0.43 0.53 0.47 0.48 0.48 0.54 0.46 0.51
MNSFS 0.50 0.50 0.23 0.45 0.31 0.48 0.47 0.34 0.44 0.40
GIN 0.54$\pm$0.05 0.55$\pm$0.05 0.63$\pm$0.22 0.59$\pm$0.03 0.57$\pm$0.07 0.56$\pm$0.03 0.56$\pm$0.07 0.70$\pm$0.20 0.56$\pm$0.04 0.61$\pm$0.06
ST-GCN 0.49$\pm$0.02 0.50$\pm$0.01 0.84$\pm$0.10 0.54$\pm$0.02 0.62$\pm$0.03 0.51$\pm$0.04 0.52$\pm$0.08 0.66$\pm$0.03 0.48$\pm$0.06 0.51$\pm$0.22
CORAL 0.59$\pm$0.06 0.61$\pm$0.05 0.56$\pm$0.05 0.61$\pm$0.06 0.59$\pm$0.05 0.55$\pm$0.03 0.57$\pm$0.02 0.57$\pm$0.04 0.60$\pm$0.05 0.57$\pm$0.03
DANN 0.59$\pm$0.03 0.59$\pm$0.03 0.60$\pm$0.08 0.60$\pm$0.03 0.59$\pm$0.04 0.55$\pm$0.05 0.54$\pm$0.04 0.70$\pm$0.23 0.59$\pm$0.04 0.60$\pm$0.09
AUFA (Ours) 0.61$\pm$0.03 0.60$\pm$0.03 0.73$\pm$0.03 0.63$\pm$0.02 0.66$\pm$0.03 0.62$\pm$0.05 0.61$\pm$0.04 0.78$\pm$0.05 0.68$\pm$0.04 0.68$\pm$0.03
V Experiments
V-A Experimental Settings
In this work, we designed 4 cross-site fMRI prediction tasks: SITE 20 $\rightarrow$ SITE 21, SITE 20 $\rightarrow$ SITE 25, SITE 20 $\rightarrow$ SITE 1, SITE 20 $\rightarrow$ SITE 9. The dataset in front of the arrow is defined as the source domain, and the dataset after the arrow is defined as the target domain. The source domain dataset contains full category labels and the target domain dataset does not have any labeling information. Considering the limited number of samples, we use all source/target domain samples to train and test the categories to which all target domain samples belong. We trained the model with 5 random initializations and recorded the mean and standard deviation results. The model was evaluated using 5 metrics, including (1) accuracy, (2) precision, (3) recall, (4) area under the receiver operating characteristic curve (AUC), and (5) F1-score.
Figure 3: Visualization of feature distributions at different sites before and after domain adaptation by t-SNE [43], where yellow ”+” denotes the source domain and green dots denote the target domains.
V-B Competition Methods
We compare the proposed AUFA with three traditional (BC-SVM, LE-SVM and MNSFS) and four deep learning methods (GIN, ST-GCN, CORAL, and DANN). Among them, BC-SVM, LE-SVM, MNSFS, GIN and ST-GCN have not undergone adaptive learning, meaning that the model trained on the source domain data is directly used to predict the target domain data. Note that CORAL, DANN and AUFA are different domain adaptation methods, and unlabeled target domain data will be used for training a specific model.
(1) BC-SVM: In this method, we construct an FCN for each subject based on PC. The betweenness centrality (BC) [44] of each node (ROI) is then extracted from the FCN to measure the number of shortest paths between that node and other nodes in the network. Finally, the corresponding FCN for each subject has a 116-dimensional feature and is classified by the SVM.
(2) LE-SVM: Similar to BC-SVM, we extract the local efficiency [45] of each node as features ($116$ dimension) based on the constructed FCN, followed by SVM for classification.
(3) MNSFS: The method extracts multiple measures for each node from the estimated FCN, including three different definitions of local clustering coefficients, four centralities (i.e., degree centrality, betweenness centrality, closeness centrality, and eigenvector centrality), as well as local efficiency. The measures belonging to each node are assigned to a group, followed by sparse group LASSO for feature selection and SVM for classification [46].
(4) GIN: It uses graph isomorphic network (GIN) [47] for FCN feature extraction. Each GIN layer aggregates the neighbor features of each node based on FCN and uses MLP to perform nonlinear transformation to achieve node update. This process can be repeated through multiple layers of the same structure to gradually fuse the global information. Finally, the output of the GIN module is average pooled to obtain the graph representation features for each subject, and classify them by the fully connected layer. The number of GIN layers is $2$, and the output dimension of each layer is $64$.
(5) ST-GCN: It models dynamic features of fMRI through spatio-temporal graph convolutional networks (ST-GCN) [48]. We input regional fMRI time series into two spatio-temporal graph convolutional units to learn the spatio-temporal representation of fMRI. Finally, the classification tasks of MDD and NC are performed through fully connected layers. The output dimension of each convolutional unit is $64$.
(6) CORAL: In this method, we use the same graph representation learning method as the proposed AUFA to extract features from each subject’s FCN. Here, we use the label of the source domain for supervised training, and apply the covariance alignment (CORAL)  [49] module to minimize the distribution differences between the source and target domain data, thereby reducing the effect of domain offset.
(7) DANN: The domain adversarial neural network (DANN) [50] is a deep learning model for domain adaptation. In DANN, a convolutional neural network is used to extract the features of the source and target domains, and the domain discriminator is used to judge which domain the data comes from. Similar to AUFA, DANN only has supervision of source domain and learns domain invariant features through adversarial training of domain discriminators and classifiers.
V-C Experimental Results
In Table II, we report the experimental results of our proposed AUFA and $7$ comparison methods in MDD identification tasks. Statistical significance analysis results are reported in Supplementary Materials. From the Table II, we can make the following observations. First, the methods employing domain adaptive strategies (i.e., CORAL, DANN, and our proposed AUFA) exhibit significant advantages when compared to learning methods without domain adaptive processing (i.e., BC-SVM, LE-SVM, MNSFS, GIN, and ST-GCN). This phenomenon reveals the limitation on the generalization performance of models trained directly from the source domain to the target domain due to the significant data heterogeneity problem among different sites. This validates the effectiveness of utilizing the maximum mean difference constraint to achieve inter-domain feature alignment. Second, deep learning models such as GIN and ST-GCN show relatively high performance in methods that do not perform domain adaptation. This is likely due to their capability to adaptively learn discriminative features that are beneficial for the classification task under label supervision. In contrast, traditional machine learning methods heavily rely on predefined features and prior knowledge. Furthermore, our AUFA model achieves excellent performance on most of the metrics in several domain adaptive methods. Compared with CORAL, which adopts the same graph representation learning in our method, AUFA achieves an average improvement of about 5% and 4% in ACC and AUC metrics on the four domain adaptation tasks. This improvement can be attributed to the augmentation-based self-optimization module integrated within AUFA, which effectively alleviates the problem of model overfitting on the source domain, resulting in beneficial optimization of the overall model performance. Finally, compared to the popular DANN method, AUFA still maintains outstanding performance with an average improvement of about 5.5% and 5% in ACC and AUC metrics across the four domain adaptation tasks. Besides the augmentation of target data by AUFA, which mitigates the risk of overfitting, another possible reason is that the convolutional neural network structure adopted by DANN may not fully recognize and focus on the importance of features from different brain regions. In contrast, AUFA leverages self-attention mechanisms to learn different weights for each brain region, thus extracting more discriminative features.
V-D Visualization of FMRI Data Distribution
We visualize the distribution differences between the features at each site before and after domain adaptation. Specifically, we flattened the upper triangular matrix of the FCNs of the source and target domain data as the features without domain adaptation, and then, the features learned by the Transformer in our AUFA framework as the features after domain adaptation. Finally, these features are embedded into a 2D space for visualization by t-SNE [43], with results displayed in Fig. 3. As can be seen in Fig. 3, the feature distributions of the two sites that have not undergone domain adaptation show a large domain offset. After domain adaptation, the distributions of the source and target domains become consistent. This indicates that our method can effectively reduce the heterogeneity of the fMRI data from different sites and thus improve the performance of the model on the target domain.
VI Discussion
VI-A Ablation Experiment
To explore the effectiveness of the AUFA framework more deeply, we designed three experiments to further analyze the effect of each sub-module: (1) Feature extraction is performed using the same backbone network as AUFA, but without any domain adaptation and augmentation operations (only $L_{C}$ is retained), i.e., AUFA-C; (2) The domain adaptation module in the AUFA framework is discarded and only $L_{C}$ and $L_{A}$ are retained, i.e., AUFA-AUG; (3) Without augmenting the data in the target domain, only $L_{C}$ and $L_{M}$ are retained, i.e., AUFA-MMD. We conduct ablation experiments on the largest sample classification task (i.e., SITE 20 $\rightarrow$ SITE 21), and the results are shown in Table III. It can be observed that our AUFA obtains the best classification results, and removing any part of the framework results in a degradation of performance. This demonstrates the effectiveness of our proposed framework, in which the maximum mean difference constraint effectively achieves inter-domain feature alignment. Additionally, the augmentation of the target domain alleviates overfitting and improves model generalizability.
TABLE III: Classification results of our AUFA and its three variants in the MDD and NC classification task for cross-domain SITE 20 $\rightarrow$ SITE 21, with the best results shown in bold.
Method Accuracy Precision Recall AUC F1-score
AUFA-C 0.54$\pm$0.03 0.59$\pm$0.03 0.58$\pm$0.04 0.58$\pm$0.01 0.58$\pm$0.02
AUFA-AUG 0.55$\pm$0.01 0.59$\pm$0.01 0.62$\pm$0.05 0.56$\pm$0.01 0.60$\pm$0.02
AUFA-MMD 0.56$\pm$0.01 0.60$\pm$0.01 0.61$\pm$0.03 0.56$\pm$0.01 0.60$\pm$0.02
AUFA (Ours) 0.63$\pm$0.03 0.66$\pm$0.03 0.73$\pm$0.06 0.62$\pm$0.04 0.69$\pm$0.03
VI-B Impact of Hyperparameters
[…]
Figure 4: Accuracy (ACC) achieved by AUFA in MDD and NC classification with different values of $\lambda_{1}$ and $\lambda_{2}$.
VI-C Discriminative Brain Regions
This work addresses the problem of data heterogeneity in multi-site fMRI while identifying the most discriminative brain regions in the task of MDD vs. NC classification. In Section IV-B1, the self-attention mechanism we employed to learn the attention matrix $A$ can efficiently quantify the interdependence of different brain regions. Given that our proposed AUFA framework is used for brain disorder detection tasks, the elements with larger values in $A$ indicate those brain region connections that play a decisive role in distinguishing healthy from diseased states. To gain deeper insight into these critical connections, we averaged the multiple attentional matrices learned for each attention head, aiming to strengthen those significant connections consistently apparent across multiple attentional perspectives. Subsequently, the top ten most discriminative brain region connections were visualized by the BrainNet Viewer [51] tool, and the results are shown in Fig. 5. As shown in Fig. 5, the identified brain regions such as the calcarine fissure (CAL), superior temporal gyrus (STG), lingual gyrus (LING), and medial and cingulum gyrus (DCG) play important roles. This is highly correlated with previous studies of MDD [52, 53, 54, 55], thus validating the effectiveness of AUFA based on rs-fMRI in identifying disease-related functional connectivity patterns. In the Supplementary Materials, we also elaborate on the influence of different source domains.
VI-D Limitations and Future Work
The current work still has several limitations. First, The proposed framework only considers resting-state fMRI data, while other modalities (e.g., structural MRI and demographic information) can provide rich complementary information about the brain from different perspectives. Employing multi-modality data to help boost learning performance, which will also be our further work. Second, this work mainly focuses on dealing with the migration problem from a single source domain to a single target domain. We can develop more advanced models to integrate and optimize information from different source domains, and explore effective adaptive strategies for complex scenarios involving multiple source domains and multiple target domains. In addition, it is interesting to generalize the model to unknown domains. In the future, we will focus on constructing and optimizing generative models for modeling data differences between source and target domains to facilitate knowledge migration and generalization, and to further improve the generalization performance of our method over unknown domains.
Figure 5: Visualization of the top ten most discriminative functional connections detected by our method in the task of MDD vs. NC classification using the BrainNet Viewer [51].
VII Conclusion
In this paper, we propose an augmentation-based unsupervised cross-domain fMRI adaptive framework for the automatic diagnosis of major depressive disorder (MDD). We aim to reduce the distributional differences of fMRI data from different sites and address the overfitting of the model in the source domain. We learn graph-level features with spatial attention via Transformer encoder and mitigate data heterogeneity between domains via the domain adaptation module. We also augment the target domain data to further improve the generalization of the model in the target domain. Experimental results on multiple sites demonstrate the effectiveness of AUFA in MDD detection. In addition, our framework can discover discriminative brain regions, providing potential imaging biomarkers for MDD analysis.